If a sausage-making machine produces 3000 sausages in h hours, how many sausages can it produce in m minutes?
Note: I can plug in numbers, but am struggling to set this problem up algebraically. I think I'm solving for s, in terms of m....
There are 5 possible answers:
a. m/50h
b.50h/m
c. mh/50
d. 50m/h
e. 50mh
I can plug in values to see which one works (d works), but would love to figure out how to solve this problem through the use of algebra...

Answer by stanbon(75887)

If a sausage-making machine produces 3000 sausages in h hours, how many sausages can it produce in m minutes?
Ans:
m minutes = (m/60) hrs.
-----
Let the number of sausages in m minutes be "x":
Solve for "x":
x/(m/60) = 3000/h
---
x = (m/60)(3000/h)
---
x = 50m/h sausages
======================
Ans: d
======================
Cheers,
Stan H.
Note: I can plug in numbers, but am struggling to set this problem up algebraically. I think I'm solving for s, in terms of m....
There are 5 possible answers:
a. m/50h
b.50h/m
c. mh/50
d. 50m/h
e. 50mh